Five years ago, John's age was half of the age he will be in 8 years. How old is he now?
Answer by stanbon(75887)

Five years ago, John's age was half of the age he will be in 8 years. How old is he now?
-------------------
Equation:
(j-5) = (1/2)(j+8)
j-5 = (j/2) + 4
(1/2)j = 9
j = 18 yrs old (John's age now)
==================================
Cheers,
Stan H.